# RoundTable: Investigating Group Decision-Making Mechanism in Multi-Agent Collaboration

### v1

@@ D.3.1 Linguistic Statistics
AVERAGE INFORMATION DIFFERENCE
@@
D.4 Dialogue Acts
@@
Figure 5: Dialogue act transition graph for different social choices in exchange economics environment.
Figure 6: Dialogue act transition graph for different social choices in recommendation system environment.
D.5 Early Stopping
Table 10: Comparison between early stopping methods in different social choices. The performance is shown with group total utility and MAE. ($\uparrow$) and ($\downarrow$) indicates better performance with higher and lower values, respectively. Experiments are based on the results of 3 agents, gpt-4o-mini setting. Results are based on 5-fold cross validation. We observe that language-based methods performed well overall.
@@ D.5 Early Stopping
Table 11: Basic statistics of different early stopping methods.
@@
Threshold Information Difference 0.17 0.17 0.17 0.20 0.18 0.17 0.18
@@
Threshold Information Difference 0.15 0.13 0.12 0.14 0.12 0.14 0.13
We share full result of early stopping experiment in Table 10, and basic statistics of early stopping methods, including early stopped round, effective ratio, and information difference threshold in Table 11. Effective ratio stands for the chance of a certain rule can be applied in the test set. Threshold is the embedding distance threshold used for early stopping.
Although linguistic feature based methods, Information Difference and Dialogue Act, outperforms other early stopping methods and baseline in exchange economy environment, the difference is small. This is due to the small room for improvement, identified between the baseline (@10) and Oracle performance.

### v2

@@ E.3 Cross-Agent Communication Analysis
AVERAGE INFORMATION DISTANCE
@@
Table 8: Heatmaps of average message length, message complexity, and information distance in each round. Analysis is made in 3 agents, gpt-4o-mini setting. Color gradient is calculated with green as maximum, red as minimum, and white as median value in each table.
The dialogue act annotation results are presented in Table 9. The dialogue act annotation results in both environments show that Inform and Request acts dominate, indicating frequent information sharing and input requests critical for task progression. The Confirm act rises sharply after round 3 in the recommendation system and more gradually in the exchange economy, showing validation is more prominent in recommendation tasks. Summarize acts increase slightly in later rounds, consolidating information, while Evaluate acts remain consistently high, reflecting ongoing assessment. Propose acts surge early in the exchange economy and peak later in the recommendation system, suggesting early proposals are vital in negotiations. Acts like Compromise, Defend, Accept, and Decline are rare, reflecting minimal adversarial behavior, while Others remain low, indicating the high quality of dialogue act definitions.
(a) Dialogue act transition graph for exchange economy.
(b) Dialogue act transition graph for recommendation system.
The most probable transition graphs in both environments are shown in Figure 22(a) and 22(b). The breakdown of voting rules are displayed in Figure 23 and 24. The transitions reveal a structured progression from requesting information to proposing and evaluating solutions, followed by resolution or negotiation. In the exchange economy, collaboration centers around a loop between Request and Propose, reflecting the cooperative decision-making process. Key transitions include Inform to Request, Confirm to Request, and Compromise to Propose, showing how agents share knowledge, confirm information, and adopt ideas. Accept transitions to End, signaling negotiation closure. In contrast, the recommendation system centers on Request, indicating a more cooperative, exploratory process with incomplete information.
@@
Figure 23: Dialogue act transition graph for different voting rules in exchange economics environment.
Figure 24: Dialogue act transition graph for different voting rules in recommendation system environment.
E.4 Early Stopping
@@
Table 10: Comparison between early stopping methods in different voting rules. The performance is shown with group total utility and MAE. ($\uparrow$) and ($\downarrow$) indicates better performance with higher and lower values, respectively. Experiments are based on the results of 3 agents, gpt-4o-mini setting. Results are based on 5-fold cross validation. We observe that language-based methods performed well overall.
@@
Threshold Information Distance 0.17 0.17 0.17 0.20 0.18 0.17 0.18
@@
Threshold Information Distance 0.15 0.13 0.12 0.14 0.12 0.14 0.13
Table 11: Basic statistics of different early stopping methods.
We share full result of the early stopping experiment in Table 10, and basic statistics of early stopping methods, including early-stopped round, effective ratio, and information distance threshold in Table 11. The effective ratio stands for the chance that a certain rule can be applied in the test set. The threshold is the embedding distance threshold used for early stopping.
Although language feature based methods, Information Distance and Dialogue Act, outperforms other early stopping methods and baseline in exchange economy environment, the difference is small. This is due to the small room for improvement, identified between baseline (@10) and Oracle performance.